# A Trifecta of Modelling Tools: Navigating COMPAS, Shark and Bayesian Inference for Binary Black Hole Model Selection

### v1

@@ 2.1.1 Common envelope evolution ($\alpha_{\rm CE}$) and Mass transfer efficiency ($f_{\rm MT}$)
Given the above, in the formation of isolated binary stars, uncertainties remain in the common envelope (CE) phase and mass transfer process, which play a key role in the final component and orbital properties.
@@
with $0\leq f_{\rm MT}\leq 1$. This fraction depends on the properties of the binary prior to mass transfer and varies over time. We use a fixed fraction prescription in COMPAS, so the fraction of mass accreted is assumed to be constant over the mass transfer phase. $f_{\rm MT}=1$ corresponds to fully conservative mass transfer, so all the mass lost from the donor is accreted. $f_{\rm MT}<1$ is non-conservative and $f_{\rm MT}=0$ is fully non-conservative. In the default setting, the mass transfer efficiency is determined by the ratio of the Kelvin–Helmholtz thermal timescale of the two stars,
@@
This is the time taken for the star’s internal energy, $E_{\rm int}$, to be radiated at it’s current luminosity, $L$. The maximum accretion rate onto a stellar accretor is,
@@
where the black hole radius, $R_{*}$, is the Schwarzschild radius. In the unstable mass transfer scenario, for simplicity, the CE phase is parameterised by COMPAS in the context of energy balance with the $\alpha_{\rm CE}$ formalism (Xu & Li, 2010). The CE efficiency, $\alpha_{\rm CE}$, describes the efficiency of converting the orbital energy into kinetic energy to unbind the common envelope. This unbinding energy is defined as,
$E_{\rm bind}=\alpha_{\rm CE}\left(-\frac{GM_{d}M_{2}}{2a_{i}}+\frac{GM_{\rm core}M_{2}}{2a_{f}}\right),$ (5)
@@
where G is the gravitational constant, $M_{d}$ is the mass of the pre-CE donor, $M_{2}$ is the mass of the secondary, $M_{\rm core}$ is the mass of the donor’s core, $a_{i}$ and $a_{f}$ are the pre- and post-CE orbital separations. Increasing $\alpha_{\rm CE}$ increases the unbinding energy, so the CE can be ejected more efficiently at wider binary separations. Too small values of $\alpha_{\rm CE}$ would require the binary to shrink in order to obtain enough unbinding energy (Tutukov & Yungelson, 1979; Webbink, 1984; Ivanova et al., 2013; Hirai & Mandel, 2022; Stevenson & Clarke, 2022). de Kool (1990) proposed an alternative version of the binding energy, introducing a new parameter, $\lambda$, to characterise the central concentration of the donor’s envelope,
$E_{\rm bind}=-\frac{GM_{1}M_{\rm env}}{\lambda a_{i}r_{L}},$ (6)
@@
where $M_{\rm env}=M_{1}-M_{\rm core}$ is the mass of the envelope and $r_{L}=R_{L}/a_{i}$ is the ratio of the Roche lobe radius and pre-CE orbital separation. The post-CE orbital separation is determined by substituting Eq 6 into Eq 5,
$\frac{a_{f}}{a_{i}}=\frac{M_{\rm core}M_{2}}{M_{1}}\frac{1}{M_{2}+2M_{\rm env}/\alpha_{\rm CE}\lambda r_{L}}.$ (7)
Due to lack of understanding and convenience in population synthesis calculations, $\alpha_{\rm CE}$ and $\lambda$ are set to constants, but should be varied depending the stellar and binary parameters (Xu & Li, 2010).
@@
Progenitors of GW sources provide an alternative path for comparison between models and observations. Zuo & Li (2014) uses the binary population synthesis code by Hurley et al. (2000) and Hurley et al. (2002) to study high-mass X-ray binaries (HMXB) and the correlation between their apparent luminosity and displacement from star clusters. A high value of $\alpha_{\rm CE}=0.8-1.0$ is preferred when comparing to observations by Kaaret et al. (2004). However, the ambiguity in the definition of the donor core mass boundary introduces a large uncertainty (almost two orders of magnitude), which translates directly to the expected value of $\alpha_{\rm CE}$. Romero-Shaw et al. (2023) uses COMPAS to study black hole high-mass X-ray binaries (BH-HMXB). They found for the case where $f_{\rm MT}\leq 0.1$ (and $f_{\rm WR}=0.1$, explained in Section 2.1.2), that $8\%$ of BH-HMXBs end up as BBHs or BHNSs. de Mink et al. (2007) attempts to fit the mass transfer efficiency fraction for massive close binaries. They find a wide spread in $f_{\rm MT}$ and conclude that a single value is not valid. Therefore, more electromagnetic observations are required to constrain modelling degeneracies of massive binaries.
@@
There have been many studies on the mass transfer and common envelope phase. Many of these have explored how varying these parameters affect the BBH merger rate, the BH mass distribution range and which binary formation channels dominate these BBH mergers (Santoliquido et al., 2021; Grichener, 2023; van Son et al., 2020, 2022a, 2022b; Marchant et al., 2021). To infer the formation channels of BBH mergers, many studies further explore the connection between the models and data (Cheng et al., 2023b). Bouffanais et al. (2021) find the preferred parameter range for the common envelope and mass transfer efficiency is $[4,7]$ and $[0.7,0.8]$ respectively. They also showed that increasing the common envelope efficiency increases the uncertainty on the local BBH merger rate density. However studies such as Bavera et al. (2021),Klencki et al. (2021) and Cui & Li (2023) find there is a preference for lower $\alpha_{\rm CE}$ ($10.2<\alpha_{\rm CE}<$ 1, $\alpha_{\rm CE}\leq 0.7$ and $\alpha_{\rm CE}=0.5,1$ respectively). Klencki et al. (2021) further questions the validity of the CE channel for producing massive BHs. They find that successful ejection of the CE is possible if the donor is a massive convective-envelope giant.
@@
We briefly mention the various studies that explore the uncertainties in parameters. But another uncertainty to note is in how these parameters are implemented in population synthesis codes. Gallegos-Garcia et al. (2021) highlights the uncertainty in stellar evolution tracks by comparing the mass transfer and common envelope treatments in the population synthesis codes MESA and COSMIC. The COSMIC models primarily form BBH mergers through the common envelope phase, while the stable mass transfer formation channel is preferred with MESA. Increasing $\alpha_{\rm CE}$ also increases the variation in the BBH merger rates between the models, due to variations in merger times. Agrawal et al. (2022a) and Romagnolo et al. (2023) have compared population synthesis codes to explore the uncertainty in radial expansion of massive stars and the variation in the remnant mass.
@@ 2.1.2 Wolf–Rayet Winds ($f_{\rm WR}$)
where $M$ is the mass of the helium star, $L$ is the luminosity, $Z$ is the metallicity, $Z_{\odot}=0.014$ (Asplund et al., 2009) and $m=0.86$ (Vink & de Koter, 2005). Wolf–Rayet stars have been studied to understand their explosion mechanisms and determine if they are progenitors of gamma ray bursts (Detmers et al., 2008). More recently, Wolf–Rayet stars have been used to study GW sources (Vink & Harries, 2017; Bogomazov et al., 2018). Decreasing $f_{\rm WR}$ leads to an increase in the BH masses (Dorozsmai & Toonen, 2022; Romero-Shaw et al., 2023).
@@
HMXBs are usually ideal candidates to study and constrain the effects of stellar mass loss in interacting binaries, as radiation driven winds are expected for giant stars. Schrøder et al. (2021) finds that the orbital evolution of HMXBs depends on the wind velocity and mass ratio. Slow winds and high mass objects results in the binary shrinking, due to drag forces. In the case where there is no drag, the binary separation widens. GWs can provide an indirect approach to study the separation of binaries, as their progenitors include HMXBs. However, Liotine et al. (2023) found that no currently observed HMXB is predicted to form a merging BBH with high probability. They use COSMIC (Breivik et al., 2020) to conclude that the BHs in the detectable HMXBs are not likely to have masses $>35M_{\odot}$, the detectable HMXBs will not merge as BBHs within the Hubble time and the distribution of redshifts, masses and metallicities of detectable HMXBs and BBHs are different. Fishbach & Kalogera (2022) argue that discrepancies in the BH masses in BBHs and HMXBs are due to GW observational selection effects.
@@
Neijssel et al. (2021) and Romero-Shaw et al. (2023) find a preference for low $f_{\rm WR}$ ($<0.2$) for HMXB systems such as Cygnus X-1 (Miller-Jones et al., 2021). Stevenson & Clarke (2022) explores the correlated impact of model parameters on the merger rate. They identified a correlation between the Wolf-Rayet factor and metallicity, as increasing both parameters increases mass loss via stellar winds. They found most models that match the current observations have $f_{\rm WR}>1$. This contradicts the previously mentioned studies due to COMPAS over-predicting the rates and masses of BBHs formed through chemically homogeneous evolution (see Section 2.2).
@@
Riley et al. (2021) studies the CHE binaries as a formation channel for various $f_{\rm WR}$ with COMPAS. They found a correlation between the delay time and $f_{\rm WR}$, as increasing mass loss leads to wider binaries. This results in the merger rate decreasing with increasing $f_{\rm WR}$. For $f_{\rm WR}=1$, the CHE channel contributes $20$ $\rm Gpc^{-3}yr^{-1}$ to the total rate of $50$ $\rm Gpc^{-3}yr^{-1}$, which is an overestimation compared to GWTC-3. Due to the short delay times, the BBH merger rate peaks at higher redshifts ($z\approx 3-4$) compared to the peak SFR.
2.1.3 Specific angular momentum loss ($f_{\gamma}$)
@@
Mass transfer from a massive donor to a lower-mass accretor adds specific angular momentum to the transferred mass, so conservation of angular momentum requires that binary’s orbit shrinks. If mass transfer is non-conservative, angular momentum, $J_{\rm orb}$, is carried away by mass lost from the system. The change in angular momentum is parameterised by $\gamma$,
@@
$=-2\frac{\dot{M}_{d}}{M_{d}}\left[1-\frac{f_{\rm MT}}{q}-(1-f_{\rm MT})\left(\gamma+\frac{1}{2}\right)\frac{1}{1+q}\right].$ (10)
Willcox et al. (2023) introduces a new parametrisation, $a_{\gamma}$,
@@
which is the effective decoupling radius between the accretor, $a_{\rm acc}$, and the L2 point, $a_{\rm L2}$. COMPAS default model assumes isotropic re-emission of matter during non-conservative stable mass transfer ($f_{\gamma}=0$). In this model, mass lost from the donor is transported near the accretor and ejected via fast isotropic winds. Hydrodynamical simulations indicate that this mass is lost between the accretor and L2 point (MacLeod et al., 2018a, b). $f_{\gamma}$ determines the point where this angular momentum is lost. If the donor star is growing rapidly with respect to its Roche Lobe, the system will lose mass through the outer Lagrangian point (L2), which causes the orbit to shrink. This corresponds to $f_{\gamma}=1$. Angular momentum loss directly from the accretor corresponds to $f_{\gamma}=0$ (Costa et al., 2023; Romero-Shaw et al., 2023).
Lu et al. (2023) and Marchant et al. (2021) found that their models predicted high mass transfer rates, which lead to mass loss through the L2 point. MacLeod et al. (2018a); MacLeod & Loeb (2020) argue that the specific angular momentum is ejected between the accretor and the L2 point. Willcox et al. (2023) uses COMPAS to study the stable mass transfer channel. They found that increasing $f_{\gamma}$ leads to a reduction in the number of systems that undergo stable mass transfer. Increasing $f_{\rm MT}$ while decreasing $f_{\gamma}$ increases the likelihood of stable mass transfer. For $f_{\gamma}\gtrsim 0.7$, nearly all systems experience common envelope evolution.
van Son et al. (2022b) studies the impact of $f_{\gamma}$ on the local BBH merger rate and the primary BH mass distribution. for $f_{\gamma}=0.75$ and $f_{\gamma}=1$, there are no BBHs formed through the stable mass transfer channel. The merger rate drops from $141\rm Gpc^{-3}yr^{-1}$ with $f_{\gamma}=0.25$ to $16\rm Gpc^{-3}yr^{-1}$ with $f_{\gamma}=0.5$, which implies the stable mass transfer contributes strongly to the merger rate for $f_{\gamma}<0.5$. As the low mass end of the primary BH mass distribution is dominated by the stable mass transfer channel, as $f_{\gamma}$ increases, the minimum primary BH mass increases.

### v2

@@ 2.1.1 Common envelope evolution ($\alpha_{\rm CE}$) and Mass transfer efficiency ($f_{\rm MT}$)
Given the above, in the formation of isolated binary stars, uncertainties remain in the CE phase and mass transfer process, which play a key role in the final component and orbital properties.
@@
with $0\leq f_{\rm MT}\leq 1$ and $\dot{M}_{a}$ and $\dot{M}_{d}$ are the mass loss rates for the accretor and donor respectively. This fraction depends on the properties of the binary prior to mass transfer and varies over time. We use a fixed fraction prescription in COMPAS, so the fraction of mass accreted is assumed to be constant over the mass transfer phase. $f_{\rm MT}=1$ corresponds to fully conservative mass transfer, so all the mass lost from the donor is accreted. $f_{\rm MT}<1$ is non-conservative and $f_{\rm MT}=0$ is fully non-conservative. We choose to explore the full range of possibilities in COMPAS. We also explore the default setting, where the mass transfer efficiency is determined by the ratio of the Kelvin–Helmholtz thermal timescale of the two stars,
@@
This is the time taken for the star’s internal energy, $E_{\rm int}$, to be radiated at its current luminosity, $L$. This ratio translates to a fixed mass ratio, $0\leq f_{\rm MT}\leq 1$, but depends in the initial mass ratio and orbital separation (Schneider et al., 2015). The maximum accretion rate onto a stellar accretor is,
@@
where the black hole radius, $R_{*}$, is the Schwarzschild radius. The Eddington accretion limit occurs when the BH accretion disk reaches a maximum luminosity beyond which radiation pressure will overcome gravity, and the infall of material outside of the object will be halted. van Son et al. (2020) show that it is possible to break this limit via stable mass transfer or CE accretion, known as super-Eddington accretion. In the unstable mass transfer scenario, for simplicity, the CE phase is parameterised by COMPAS in the context of energy balance with the $\alpha_{\rm CE}$ formalism (Xu & Li, 2010) (see also Ivanova et al., 2013, and references therein). The CE efficiency, $\alpha_{\rm CE}$, describes the efficiency of converting the orbital energy into kinetic energy to unbind the CE. This unbinding energy is defined as,
$E_{\rm bind}=\alpha_{\rm CE}\left(\frac{GM_{d}M_{a}}{2a_{i}}-\frac{GM_{\rm d,core}M_{\rm a,core}}{2a_{f}}\right),$ (5)
@@
where G is the gravitational constant, $M_{d}$ is the mass of the pre-CE donor, $M_{a}$ is the mass of the accretor, $M_{\rm d,core}$ is the mass of the donor’s core, $a_{i}$ and $a_{f}$ are the pre- and post-CE orbital separations respectively. Increasing $\alpha_{\rm CE}$ increases the unbinding energy, so the CE can be ejected more efficiently at wider binary separations. Too small values of $\alpha_{\rm CE}$ would require the binary to shrink in order to obtain enough unbinding energy (van den Heuvel, 1976; Tutukov & Yungelson, 1979; Webbink, 1984; Ivanova et al., 2013; Marchant et al., 2021; Hirai & Mandel, 2022; Stevenson & Clarke, 2022). de Kool (1990) proposed an alternative version to the $\alpha$-formalism for the binding energy, introducing a new parameter, $\lambda$, to characterise the central concentration of the donor’s envelope,
$E_{\rm bind}=-\frac{GM_{d}M_{\rm env}}{\lambda a_{i}r_{L}},$ (6)
@@
where $M_{\rm env}=M_{\textrm{d}}-M_{\rm d,core}$ is the mass of the envelope and $r_{L}=R_{L}/a_{i}$ is the ratio of the Roche lobe radius and pre-CE orbital separation. The post-CE orbital separation is determined by substituting Eq. 6 into Eq. 5,
$\frac{a_{f}}{a_{i}}=\frac{M_{\rm d,core}M_{a}}{M_{d}}\frac{1}{M_{a}+2M_{\rm env}/\alpha_{\rm CE}\lambda r_{L}}.$ (7)
Due to lack of understanding and convenience in population synthesis calculations, $\alpha_{\rm CE}$ and $\lambda$ are set to constants, but should be varied depending the stellar and binary parameters (Xu & Li, 2010). For $0\leq\alpha_{\rm CE}\leq 1$, the only source of energy to unbind the CE is the orbital energy. However, various hydrodynamical simulations over the years have shown that the orbital energy alone might not be sufficient for envelope removal (Ivanova et al., 2013; Röpke & De Marco, 2023). Additional energy sources are required ($\alpha_{\rm CE}>1$) such as recombination energy from evolved giant star envelopes (Ivanova et al., 2015; Lau et al., 2022) or the giant stars can launch jets as they accrete mass via an accretion disk (Ohlmann et al., 2016; Moreno Méndez et al., 2017). In this paper, we explore $0.1\leq\alpha_{\rm CE}\leq 100$. Although our upper limit seems extreme, it should be noted that Hirai & Mandel (2022) implements a two-step CE formalism which is equivalent to $\alpha_{\rm CE}>10$ in most cases (see their Fig. 5). This emphasises the need for additional sources of energy, but it may also suggest that the current $\alpha_{\rm CE}$-formalism in COMPAS is wrong. As such, we choose to include $\alpha_{\rm CE}=100$ as an exploratory parameter. See further discussion of this in Section 2.4.
@@
Progenitors of GW sources provide an alternative path for comparison between models and observations. Zuo & Li (2014) uses the binary population synthesis code by Hurley et al. (2000) and Hurley et al. (2002) to study high-mass X-ray binaries (HMXB) and the correlation between their apparent luminosity and displacement from star clusters. A high value of $\alpha_{\rm CE}=0.8-1.0$ is preferred when comparing to observations by Kaaret et al. (2004). However, the ambiguity in the definition of the donor core mass boundary introduces a large uncertainty (almost two orders of magnitude), which translates directly to the expected value of $\alpha_{\rm CE}$. Romero-Shaw et al. (2023) uses COMPAS to study black hole high-mass X-ray binaries (BH-HMXB). They find for the case where $f_{\rm MT}\leq 0.1$ (and $f_{\rm WR}=0.1$, explained in Section 2.1.2), that $8\%$ of BH-HMXBs end up as BBHs or BHNSs. de Mink et al. (2007) attempts to fit the mass transfer efficiency fraction for massive close binaries. They find a wide spread in $f_{\rm MT}$ and conclude that a single value is not valid. Therefore, more electromagnetic observations are required to constrain modelling degeneracies of massive binaries.
@@
There have been many studies on the mass transfer and CE phase. Many of these have explored how varying these parameters affect the BBH merger rate, the BH mass distribution range and which binary formation channels dominate these BBH mergers (van Son et al., 2020, 2022a, 2022b). Many of these studies also find that additional energy sources are necessary to produce BBH mergers and other double compact objects (DCO) (Santoliquido et al., 2021; Zevin et al., 2021a; Marchant et al., 2021; Grichener, 2023; Deng et al., 2024). To infer the formation channels of BBH mergers, many studies further explore the connection between the models and data (Cheng et al., 2023). Bouffanais et al. (2021) find the preferred parameter range for the CE and mass transfer efficiency is $[4,7]$ and $[0.7,0.8]$ respectively. They also showed that increasing the CE efficiency increases the uncertainty on the local BBH merger rate density. Fragos et al. (2019) find that the final orbital separation for their BNS’s to merger within a Hubble time translates to $\alpha_{\rm CE}\sim 5$. Broekgaarden & Berger (2021) and García et al. (2021) also find a preference for $\alpha_{\rm CE}>1$ for producing merger rates compared to the observed rates. However studies such as Bavera et al. (2021),Klencki et al. (2021) and Cui & Li (2023) find there is a preference for lower $\alpha_{\rm CE}$ ($10.2<\alpha_{\rm CE}<$ 1, $\alpha_{\rm CE}\leq 0.7$ and $\alpha_{\rm CE}=0.5,1$ respectively). Klencki et al. (2021) further questions the validity of the CE channel for producing massive BHs. They find that successful ejection of the CE is possible if the donor is a massive convective-envelope giant.
@@
We briefly mention the various studies that explore the uncertainties in parameters. But another uncertainty to note is in how these parameters are implemented in population synthesis codes. Gallegos-Garcia et al. (2021) highlights the uncertainty in stellar evolution tracks by comparing the mass transfer and CE treatments in the detailed stellar evolution code, MESA (Paxton et al., 2011; Paxton et al., 2013, 2015, 2018, 2019), and rapid binary population synthesis code, COSMIC (Breivik et al., 2020). The COSMIC models primarily form BBH mergers through the CE phase, while the stable mass transfer formation channel is preferred with MESA. Increasing $\alpha_{\rm CE}$ also increases the variation in the BBH merger rates between the models, due to variations in merger times. Agrawal et al. (2022a) and Romagnolo et al. (2023) have compared population synthesis codes to explore the uncertainty in radial expansion of massive stars and the variation in the remnant mass.
@@ 2.1.2 Wolf–Rayet Winds ($f_{\rm WR}$)
where $M$ is the mass of the helium star, $L$ is the luminosity, $Z$ is the metallicity, $Z_{\odot}=0.014$ (Asplund et al., 2009) and $m=0.86$ (Vink & de Koter, 2005). Wolf–Rayet stars have been studied to understand their explosion mechanisms and determine if they are progenitors of gamma ray bursts (Detmers et al., 2008). More recently, Wolf–Rayet stars have been used to study GW sources (Vink & Harries, 2017; Bogomazov et al., 2018). Decreasing $f_{\rm WR}$ leads to an increase in the BH masses (Dorozsmai & Toonen, 2022; Romero-Shaw et al., 2023). In COMPAS, we choose the Wolf–Rayet multiplier range provided in Stevenson & Clarke (2022), which accounts for the range of uncertainties in Wolf–Rayet mass-loss rates.
@@
HMXBs are usually ideal candidates to study and constrain the effects of stellar mass loss in interacting binaries, as radiation driven winds are expected for giant stars. Schrøder et al. (2021) finds that the orbital evolution of HMXBs depends on the wind velocity and mass ratio. Slow winds and high mass objects results in the binary shrinking, due to drag forces. In the case where there is no drag, the binary separation widens. GWs can provide an indirect approach to study the separation of binaries, as their progenitors include HMXBs. However, Liotine et al. (2023) find that no currently observed HMXB is predicted to form a merging BBH with high probability. They use COSMIC (Breivik et al., 2020) to conclude that the BHs in the detectable HMXBs are not likely to have masses $>35M_{\odot}$, the detectable HMXBs will not merge as BBHs within the Hubble time and the distribution of redshifts, masses and metallicities of detectable HMXBs and BBHs are different. Fishbach & Kalogera (2022) argue that discrepancies in the BH masses in BBHs and HMXBs are due to GW observational selection effects.
@@
Neijssel et al. (2021) and Romero-Shaw et al. (2023) find a preference for low $f_{\rm WR}$ ($<0.2$) for HMXB systems such as Cygnus X-1 (Miller-Jones et al., 2021). Stevenson & Clarke (2022) explores the correlated impact of model parameters on the merger rate. They identified a correlation between the Wolf-Rayet factor and metallicity, as increasing both parameters increases mass loss via stellar winds. They find most models that match the current observations have $f_{\rm WR}>1$. This contradicts the previously mentioned studies due to COMPAS over-predicting the rates and masses of BBHs formed through chemically homogeneous evolution (see Section 2.2).
@@
Riley et al. (2021) studies the CHE binaries as a formation channel for various $f_{\rm WR}$ with COMPAS. They find a correlation between the delay time and $f_{\rm WR}$, as increasing mass loss leads to wider binaries. This results in the merger rate decreasing with increasing $f_{\rm WR}$. For $f_{\rm WR}=1$, the CHE channel contributes $20$ $\rm Gpc^{-3}yr^{-1}$ to the total rate of $50$ $\rm Gpc^{-3}yr^{-1}$, which is an overestimation compared to GWTC-3. Due to the short delay times, the BBH merger rate peaks at higher redshifts ($z\approx 3-4$) compared to the peak SFR.
2.1.3 Specific AM loss ($f_{\gamma}$)
@@
Mass transfer from a massive donor to a lower-mass accretor adds specific AM to the transferred mass, so conservation of AM requires that binary’s orbit shrinks. If mass transfer is non-conservative, angular momentum, $J_{\rm orb}$, is carried away by mass lost from the system. The change in AM is parameterised by $\gamma$,
@@
$=-2\frac{\dot{M}_{d}}{M_{d}}\left[1-\frac{f_{\rm MT}}{q}-(1-f_{\rm MT})\left(\gamma+\frac{1}{2}\right)\frac{1}{1+q}\right],$ (10)
where $q=M_{a}/M_{d}$ is the mass ratio. Willcox et al. (2023) introduces a new parametrisation, $a_{\gamma}$,
@@
which is the effective decoupling radius between the accretor, $a_{\rm acc}$, and the L2 point, $a_{\rm L2}$. COMPAS default model assumes isotropic re-emission of matter during non-conservative stable mass transfer ($f_{\gamma}=0$). In this model, mass lost from the donor is transported near the accretor and ejected via fast isotropic winds. Hydrodynamical simulations indicate that this mass is lost between the accretor and L2 point (MacLeod et al., 2018a, b). $f_{\gamma}$ determines the point where this AM is lost. If the donor star is growing rapidly with respect to its Roche Lobe, the system will lose mass through the outer Lagrangian point (L2), which causes the orbit to shrink. This corresponds to $f_{\gamma}=1$. AM loss directly from the accretor corresponds to $f_{\gamma}=0$ (Costa et al., 2023; Romero-Shaw et al., 2023). We consider $0\leq f_{\gamma}\leq 1$. It is possible for cases outside of this range such as intense feedback radiation from the accretor for mass loss at or near the the inner Lagrangian point, L1, ($f_{\gamma}<0$) or magnetic fields that keep material in co-rotation beyond the L2 point ($f_{\gamma}\geq 1$) (MacLeod & Loeb, 2020; Willcox et al., 2023).
Lu et al. (2023) and Marchant et al. (2021) find that their models predicted high mass transfer rates, which lead to mass loss through the L2 point. MacLeod et al. (2018a); MacLeod & Loeb (2020) argue that the specific AM is ejected between the accretor and the L2 point. Willcox et al. (2023) uses COMPAS to study the stable mass transfer channel. They find that increasing $f_{\gamma}$ leads to a reduction in the number of systems that undergo stable mass transfer. Increasing $f_{\rm MT}$ while decreasing $f_{\gamma}$ increases the likelihood of stable mass transfer. For $f_{\gamma}\gtrsim 0.7$, nearly all systems experience CE evolution.